Zhu Manufacturing is considering the introduction of a family of new products. Long-term demand for the product group is somewhat predictable, so the manufacturer must be concerned with the risk of choosing a process that is inappropriate. Faye Zhu is VP of operations. She can choose among batch manufacturing or custom manufacturing, or she can invest in group technology. Faye won’t be able to forecast demand accurately until after he makes the process choice. Demand will be classified into four compartments: poor, fair, good, and excellent. The table below indicates the payoffs (profits) associated with each process/demand combination, as well as the probabilities of each long-term demand level.

  Poor Fair Good Excellent
Probability .1 .4 .3 .2
Batch -$ 200,000 $1,000,000 $1,200,000 $1,300,000
Custom $ 100,000 $ 300,000 $ 700,000 $ 800,000
Group Tech. -$ 1,000,000 -$ 500,000 $ 500,000 $2,000,000

A) Based on expected value, what choice offers the greatest gain?
B) What would Faye Zhu be willing to pay for a forecast that would accurately determine the level of demand in the future?
Answer #1

(A) Expected Values

Batch- 0.1*(-200000)+0.4*1000000+0.3*1200000+0.2*1300000=1000000

Custom-0.1*100000+0.4*300000+0.3*700000+0.2*800000=500000

Group Tech-0.1*(-1000000)+0.4*(-500000)+0.3*500000+0.2*2000000=250000

Based on the expected values the best choice will be Batch Process.

(B) Expected value under Certainty =0.1*100000+0.4*1000000+0.3*1200000+0.2*2000000=1170000

Faye Zhu be willing to pay for a forecast that would accurately determine the level of demand in the future = 1170000 - 1000000 =170000